shift summary for owl, 12/18/99

Ed Blucher (blucher@hep.uchicago.edu)
Sat, 18 Dec 1999 07:33:17 -0600


Shift Summary for Owl Shift, 12/18/99

On Shift: P. Gouffon, E. Blucher

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...):0

Accounting of lost time:
power failures

========================================================================
Overview of shift:
Darkness following power failures. MCR estimates at least 3 days
for recovery.

========================================================================
Description of Shift:

23:35: the ComEd line to the main substation failed (with a big boom).

23:45 We discover that we don't have a single working flashlight. We convince
RickK to leave us a flashlight from his car.

0:10 Turned off all CAMAC crates, pipeline, CsI HV, Fastbus, DC receivers,
DC HV. Also, FASTBUS alarms disabled.

1:45 MCR informs us that power will be restored temporarily by connecting the
entire site to the CR substation. Another power outage will occur when
the ComEd repair is completed (maybe on Saturday).

2:10 Power back on.

2:45 Page Cary K for vacuum system. No response.

3:10 Large Bryair heater is off. We page duty mech. to reset unit.

3:40 We are unable to communicate with Bryair from PC. We have unplugged
the autodialer. It should be plugged back in once PLC alarm is
cleared.

3:45 Another power failure. The main line going to the CR substation
failed. ComEd is investigating (don't worry about Y2K)

6:45 Power returns. MCR doesn't know whether another power outage will be
required to complete repair.

6:50 DC2 gas flow is very low.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

We can't control Bryair from DMACS terminal (contact Rich Schmitt)
- autodialer should be plugged back in once this is fixed.
Pump down vacuum tank. (Cary K pager: 722-3408)
Check DC2 gas flow.
Flashlight bulbs.